# Iterative Trajectory Exploration for Multimodal Agents

### v1

@@ 4.7 Visualization
We visualize the preference data generated by the online self-exploration framework, as shown in Figure 4. The verifier can successfully choose solutions that lead to correct intermediate results. In step 1, the solution that produces the correct brand and basic information about the smartphone is selected as the preferred data. In step 2, the solution that searches for contents most relevant to the task is selected as the preferred data. It compares the smartphone with mid-range competitors, while the rest compares the smartphone with a special one.
We visualized the task-solving procedure of the SPORT Agent compared with the SFT Agent (T3 agent), as shown in Figure 5. The SPORT Agent after the step-wise preference tuning can well solve the issues of code hallucination and tool error. For example, in case (a), the SPORT Agent correctly sends the image path to the tool. In case (b), the SPORT Agent utilizes tools to solve the task, instead of producing an answer via hallucination.
@@ 5 Conclusion
In this paper, we have presented an online self-exploration framework for multimodal agents, through which the agents can learn via automatic interaction with new environments without accessing any expert data. Based on this framework, we have presented a step-wise optimization for refining trajectories (SPORT), which can produce in-distribution preference data in complex environments. Given proper prompts, the proposed SPORT method can generate diverse multimodal tasks and make good verification of agent solutions aligned with humans. Experiments on two challenging benchmarks: GTA and GAIA, show that the proposed SPORT method achieves significant improvements on multimodal agents, demonstrating the effectiveness.
Limitation. The verifier plays an important role in the current SPORT method. However, it heavily relies on human-designed rules and prompts, causing inferior generalization for some outlines. In the future, we will explore the self-exploration techniques for the verifier, that is, learning to verify, through which the verifier can adapt to new environments with the controller together. Furthermore, we will explore the theoretical guarantee for the verifier, allowing it to scale to open settings.
Impact Statement
This paper presents work whose goal is to advance the field of AI agents. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Comparison with Existing Sampling Frameworks
Our method constructs step-level preference data for complex multimodal tasks. To contextualize its contributions, we compare against representative reinforcement-learning–based sampling frameworks along three dimensions: (1) Task Domain, (2) Collection Granularity, and (3) Annotation Format. Table 8 summarizes this comparison.
@@ Appendix A Comparison with Existing Sampling Frameworks
WebRL (Qi et al., 2024) GUI control Trajectory-Level Finetune a reward model
PAE (Zhou et al., 2024a) GUI control Trajectory-Level Use a pre-trained model
ETO (Song et al., 2024) GUI control & Embodied AI Trajectory-Level Expert labels for comparisons
DMPO (Shi et al., 2024) GUI control & Embodied AI Trajectory-Level Finetune a reward model
DigiRL (Bai et al., 2024) GUI control Step-Level Finetune a reward model
TP-LLAMA (Chen et al., 2024b) API calling Step-Level Use expert data
IPR (Xiong et al., 2024b) GUI control & Embodied AI Step-Level Use expert data
StepAgent-Inverse (Deng et al., 2024c) GUI control & Embodied AI Step-Level Use expert data
@@
    Low sampling efficiency. Generating high-quality trajectories for multimodal tasks requires expensive tool usage (e.g. LLM calls, web searches), making naive sampling impractical.
@@
Appendix B Prompt of the Step Verifier
@@
To evaluate the quality of intermediate steps taken by an agent during task execution, we design a step verifier consisting of two key components: a system prompt and a user prompt. The system prompt (Figure 6) provides the verifier model with detailed instructions for evaluating multiple candidate steps (‘CURRENT_STEP‘) based on their coherence, logical progression, and effectiveness in advancing the task. It guides the model to consider contextual alignment with the prior step, tool usage, hallucination, and content relevance. The verifier is required to select the best step and justify its decision in a structured JSON format.
@@
The following are the given task, results of previous steps, and result of the current step. TASK: <task> Step Sets: <step_set> # Format of step_set: # { #       ‘PREVIOUS_RESULT‘: <The prior results obtained by the agent.> #       ‘CURRENT_STEP‘: <The agent’s output, containing a ‘thought‘ and ‘code‘ intended to complete the task based on the observation.> #       ‘CURRENT_RESULT‘: <The result or state produced by executing the ‘CURRENT_STEP‘.> # } Now, you need to determine the best of the current steps based on the above information.
@@
Appendix C User Study Interface
C.1 Preference Alignment Study
Figure 8 (a) presents the web interface used to evaluate how well our automated verifier’s preferences align with those of human judges. In each trial, participants were shown a single task case along with a collection of candidate next-step actions (each consisting of a brief “Thought” description and an optional code snippet or tool invocation). These options were the same ones ranked by our verifier, but presented in random order to prevent positional bias.
@@ C.1 Preference Alignment Study
By comparing the human-selected option against the top choice of the verifier, we compute an agreement rate for each model and task type. High agreement indicates that the verifier captures human judgments of step quality; lower agreement reveals areas where the verifier’s ranking diverges from human intuition. These results were aggregated over 50 cases per participant, enabling both per-case analysis and overall statistics on human–verifier alignment.
C.2 Data quality
Figure 8 (b) illustrates the web interface employed in our user study. For each case, participants proceeded through two consecutive scoring phases:
  1. 1.
@@ C.2 Data quality
subfigureInterface for the user study on verifier performance. Users select their perfered steps.
subfigureInterface for the user study on task execution and trajectory. Users evaluate the quality of tasks and trajectories.

### v4

@@ 4.7 Visualization
We visualize the preference data generated by the online self-exploration framework, as shown in Figure˜4. The verifier can successfully choose actions that lead to correct intermediate results. In step 1, the action that produces the correct brand and basic information about the smartphone is selected as the preferred data. In step 2, the action that searches for content most relevant to the task is selected as the preferred data. It compares the smartphone with mid-range competitors, while the rest compares the smartphone with a special one.
We visualized the task-solving procedure of the SPORT Agent compared with the SFT Agent (T3 agent), as shown in Figure˜5. The SPORT Agent after the step-wise preference tuning can well solve the issues of code hallucination and tool error. For example, in case (a), the SPORT Agent correctly selects the web search tool while the SFT Agent uses the wrong tool with an incorrect image path. In case (b), the SPORT Agent utilizes tools to solve the task, while the SFT Agent produces an answer via hallucination.
@@ 5 Conclusion
In this paper, we have presented an online self-exploration framework for multimodal agents, through which the agents can learn via automatic interaction with new environments without accessing any annotations. Based on this framework, we have presented a step-wise optimization for refining trajectories (SPORT), which can produce in-distribution preference data in complex environments. Given proper prompts, the proposed SPORT method can generate diverse multimodal tasks and provide good verification of agent actions aligned with humans. Experiments on two challenging benchmarks, GTA and GAIA, show that the proposed SPORT method achieves significant improvements on multimodal agents, demonstrating its effectiveness.
Limitations
The verifier plays an important role in the current SPORT method. However, it heavily relies on human-designed rules and prompts, causing inferior generalization for some outliers. In the future, we will explore the self-exploration techniques for the verifier, that is, learning to verify, through which the verifier can adapt to new environments with the controller together. Furthermore, we will explore the theoretical guarantee for the verifier, allowing it to scale to open settings.
Acknowledgements
This work was supported by the Natural Science Foundation of China (NSFC) under Grants No. 62406009, No. 62172041 and No. 62176021, Shenzhen Science and Technology Program under Grant No. JCYJ20241202130548062, and Natural Science Foundation of Shenzhen under Grant No. JCYJ20230807142703006.
Appendix A Broader Impacts
SPORT’s ability to autonomously explore tool usage and refine behavior via preference feedback promises to lower the barrier to creating versatile multimodal agents, enabling researchers and practitioners—even those with limited resources—to build domain‐adapted systems for tasks such as document analysis, scientific data interpretation, and accessible educational or healthcare interfaces. By reducing reliance on costly human annotation and manual rule‐crafting, SPORT can accelerate innovation across diverse fields, fostering more inclusive and scalable AI solutions.
At the same time, increased agent autonomy raises risks of unintended bias, error propagation, and resource inefficiency. SPORT’s verifier, grounded in heuristic rules and model judgments, may inadvertently reinforce spurious behaviors or hallucinations, particularly in high‐stakes domains like legal or medical analysis. We therefore advocate for transparent auditing, human‐in‐the‐loop oversight, and careful management of exploration budgets to ensure responsible deployment.
Appendix B Comparison with Existing Sampling Frameworks
Our method constructs step‐level preference data for complex multimodal tasks. To contextualize its contributions, we compare against representative reinforcement‐learning–based sampling frameworks along three dimensions: (1) Task Domain, (2) Collection Granularity, and (3) Annotation Format. Table 8 summarizes this comparison.
@@ Appendix B Comparison with Existing Sampling Frameworks
WebRL (qi2024webrl, ) GUI control Trajectory-Level Finetune a reward model
PAE (zhou2024proposer, ) GUI control Trajectory-Level Use a pre-trained model
ETO (song2024trial, ) GUI control & Embodied AI Trajectory-Level Expert labels for comparisons
DMPO (shi2024direct, ) GUI control & Embodied AI Trajectory-Level Finetune a reward model
DigiRL (bai2024digirl, ) GUI control Step-Level Finetune a reward model
TP-LLAMA (chen2024advancing, ) API calling Step-Level Use expert data
IPR (xiong2024watch, ) GUI control & Embodied AI Step-Level Use expert data
StepAgent-Inverse (deng2024novice, ) GUI control & Embodied AI Step-Level Use expert data
@@
    Low sampling efficiency. Generating high‐quality trajectories for multimodal tasks requires expensive tool usage (e.g. LLM calls, web searches), making naive sampling impractical.
@@
Appendix C Computational Efficiency Analysis
We provide a comprehensive comparison of compute time and GPU usage between the baseline method (MAT) and our SPORT framework. Table 9 summarizes the estimated compute time and GPU cost for both methods, measured in GPU hours (GPUh) using a single A100 80GB GPU.
Data Generation Efficiency. The computational cost of Task Synthesis, Step Sampling, and Tool Calling is closely tied to the data scale. MAT synthesizes and samples steps for 20K tasks and 20K trajectories, resulting in high overhead across all three components. In contrast, SPORT performs Task Synthesis on only 2K tasks and constructs 16K preference pairs for tuning, leading to substantially reduced compute time in these stages. Specifically, SPORT reduces Task Synthesis time from 29.13h to 3.61h, Step Sampling from 69.92h to 4.53h, and Tool Calling from 58.93h to 9.86h.
Step Verification. SPORT introduces an additional Step Verification process to obtain step-level preferences, which incurs 3.36 GPU hours. While this stage adds computational cost, it remains acceptable relative to the total runtime due to the smaller data volume (16K preference pairs vs. 20K trajectories in MAT).
Training Efficiency. Model training with 4× A100 80GB GPUs shows comparable efficiency between both methods, with SPORT requiring 15.20h (60.80 GPUh) compared to MAT’s 15.77h (63.08 GPUh). This marginal difference demonstrates that SPORT’s efficiency gains primarily stem from the data generation stage rather than the training phase.
Overall Cost Reduction. As shown in Table 9, SPORT achieves a total compute time of 36.56h compared to MAT’s 173.75h, representing a 4.75× speedup. In terms of GPU cost, SPORT requires 82.16 GPUh versus MAT’s 122.01 GPUh, resulting in a 32.7% cost reduction. These substantial improvements in computational efficiency, combined with SPORT’s superior performance, validate the effectiveness of our preference-based learning approach.
Table 9: Estimated compute time and GPU cost comparison between MAT and SPORT. GPU hours (GPUh) are measured as usage of a single A100 80GB GPU for one hour. The estimates for MAT are derived from its reported tool invocation frequency, combined with empirical measurements of the GPT-4o-mini API and tool execution latency on our hardware.
Component MAT Time (h) SPORT Time (h) MAT Cost (GPUh) SPORT Cost (GPUh)
Data Generation
   Task Synthesis 29.13 3.61 0 (GPT-4o-mini API $\sim$\$500) 3.61 (Qwen2-VL-7B)
   Step Sampling 69.92 4.53 0 (GPT-4o-mini API $\sim$\$1500) 4.53 (Tuned-Qwen2-VL-7B)
   Step Verification 0 3.36 0 3.36 (Qwen2.5-7B)
   Tool Calling 58.93 9.86 58.93 9.86
Training
[…]
Total 173.75 36.56 122.01 82.16
Appendix D Task Generation
Following MAT (gao2024multimodalagenttuningbuilding, ), we first use an LLM to generate queries, and then generate both the file content and file type based on the query. Depending on the file type, we adopt different strategies for file generation:
  • For image files, we retrieve relevant images from a large image dataset (chen2024sharegpt4v, ) based on the generated file content.
  • For non-image files, .PDF, .XLSX, .DOCX, or .MP3, etc, we employ the LLM to write Python code that calls relevant libraries to convert the file content into the corresponding file format.
After generating the files, we adopt a two-step query-file verification process—Revision and Filtering—to ensure the quality of each task (i.e., query-file pair).
In the Revision step, both the query and the corresponding file are fed into a vision-language model (VLM), which is instructed to revise the query to better align with the file if necessary. For image data, we directly input the image into the VLM; for non-image data, we input the file content instead.
In the Filtering step, the VLM is no longer allowed to modify the query. Instead, it is asked to assess whether the task meets a predefined quality threshold. Only tasks that pass this quality check are retained, while all others are discarded.
D.1 Model Comparison on Task Generation
We conducted a comparative analysis of task quality between open-source and closed-source models. For this evaluation, we employed Qwen-2VL-7B (open-source) and GPT-4o-mini (closed-source) to generate 200 tasks each under identical system prompts. The resulting 400 tasks were subsequently randomized and anonymized to eliminate source bias. We recruited 20 human evaluators, with each participant assessing 20 tasks according to two key metrics: naturalness and reasonableness, rated on a 10-point scale (higher scores indicating superior quality). As demonstrated in Table˜10, the tasks produced by both models achieved comparable quality ratings, providing compelling evidence that open-source models possess sufficient capability for high-quality task generation in this domain.
Table 10: User study for open-source vs. close-source models generated tasks. Scores are scaled from 11 to $10$ and a higher score denotes better quality.
Model Task Naturalness Task Reasonableness
GPT-4o-mini 8.71 8.37
QWen-2VL-7B 8.75 8.35
Appendix E Error bar for the main results
We conduct each experiment five times and report the performance variance, as shown in Table˜11. The results demonstrate that our improvements are statistically significant compared to the observed variances.
Table 11: Performance with variance on the GTA benchmark.
Method Controller AnsAcc ToolAcc CodeExec
T3-Agent MAT Tuned Qwen2-VL-7B 53.85 64.63 84.32
SPORT Agent (Ours) Tuned Qwen2-VL-7B 60.26 ± 1.51 72.41 ± 1.11 91.87 ± 1.41
Appendix F System Prompts
We referenced the task generation prompts from MAT(gao2024multimodalagenttuningbuilding, ) for our implementation.
F.1 Prompt for Query Generation
The prompt for query generation is shown in fig.˜6.
You are tasked with generating user queries that will prompt an agent to call various tools (only use the tool listed in our toolset), including internet search capabilities, to solve real-world, practical problems. The problems should be natural, varied, and challenging, requiring the agent to reason across different domains. Ensure that the problems span a range of practical scenarios. Our toolset: TOOL_SET I will now provide examples, along with the tools. Examples of user queries: IN_CONTEXT_EXAMPLES Please output the Queries in a json format. Make sure that the queries share a similar style to the in-context examples. The output template is: $\{$ "query": "What is the weather today?", <The user query to the agent.> "tools": ["tool1", "tool2",…], <A list consisting of the tool names related to the query.>       }\},
Figure 6: Prompt for query generation.
F.2 Prompt for File Generation
The prompt for file content generation is shown in fig.˜7 and fig.˜8, and the prompt for file code generation is shown in  fig.˜9 and  fig.˜10.
You are a smart reasoner that can restore a query_solving scene between a human and an agent. Human gives a complex query and several images to the agent, and then the agent answers the query by searching on the Internet and applying tools to the images with step-by-step reasoning. Now, you will be given the query with suggested tools, I suggest you analyze the needed information to solve the query, and divide the information into two groups: searching from the Internet and extracting from the images using tools. Based on the information from the images, you need to further infer the content of these images, through which the agent could correctly solve the query. Our toolset: TOOL_SET Output MUST use the following json template. { "information": <Needed information to solve the query. For the query including creating/generating images, the information should NOT be the description of the described image.> "information from the Internet": <Information from the Internet inferences based on the given query and suggested tools. Determine which information is suitable to be obtained from the Internet. Or say no information is required from the Internet.> "information from images": <Information extracted from given images based on the suggested tools to solve the query. It should be several sentences, including information extracted from the images using tools. Determine which information is suitable to be obtained from the images, and using which tools. Do not generate image_content for the query including generating/creating an image. Or say no information is required from the images.> "file": { "image_numbers": <set an int number, the number is depended on needed information from images>, "image_content": { "image_1": <The image content should be a natural language, describing the content of the first image relevant to the query. The content should be concrete, such as concrete number, concrete name. The content should match the query and the above images.> … <if you think the query needs more than 1 image, please output image content like ’image_2’.> } } }
Figure 7: System prompt for the file content generation.
Now given the query: QUERY, firstly analyze the needed information to solve the query and divide the information into two groups: searching from the Internet or extracting from images using tools. Then for information from images, imagine possible answers for each information (it should be concrete answers instead of descriptions). Finally, output the json for the inferenced information and the content of images.
Figure 8: User prompt for the file content generation.
You are a helpful assistant and can generate a <file type placeholder> file by writing Python code. You will be given a description of the content of the file. You need to first largely extend the content, and then write Python code to generate a <file type placeholder> file. GUARANTEE that the provided content is in the file. The output Python code MUST use the following template. """ ## extention start Extened content: <here is the extented content> ## extention end ## code start <here is the Python code to generate a <file type placeholder> file> ## code end """
Figure 9: User prompt for the non-image file generation.
Now, given the following content: <file content>, first largely extend the content, and output a code to generate a <file type placeholder> file, where the file name is <file name> and the file will be saved in <save path>.
Figure 10: User prompt for the non-image file content generation.
F.3 Prompt for Query-file Filter
The prompt for the query-file filter is shown in fig.˜11 and fig.˜12.
You are a helpful assistant that is given a query and several images. You need to check whether the images are relevant to the query. The query and images are used to evaluate the perception ability, reasoning ability, and information search ability of an AI agent. The agent solves the query by searching for information on the Web and extracting information from the images. In some cases, based on the given images, the agent could not solve the query, even though it searched for information from the Web (e.g., some specific knowledge). You need to pick up these bad cases. The agent can call the following tools to solve the query. TOOL_SET . Thus, the images should follow these requirements. 1. Relevance: The depicted scenarios or objects in images should be relevant to the query. The images should contain scenarios or objects that are mentioned in the images. 2. Usefulness: The image should contain necessary information to address the query, such as some specific details that cannot be obtained from the Web. 3. Some queries require the agent to search for knowledge from the Web and combine the information in the image to solve the queries. Thus, in some cases, the images do not contain all the information to solve the query, but the missed information could be searched on the Web. These cases should be regarded as correct cases. The output MUST use the following json template to check the images. {    "information_for_query": <Required information to solve the query.>, "useful_information_in_image": <Useful information that can be extracted from images to solve the query>, "missed_information_in_images": <Missed information that is necessary to solve the query but does not exist in the images.>, "missed_information_web_search": <You need to justify whether the missed information could be searched from the Web, using your rich experience in surfing the Internet.> , "missed_information_obtained": <You need to justify whether the missed information could be obtained via computing or reasoning based on information extracted from the images or searched from the Web.>, "thought": <Now, you need to determine whether the images can solve the query. If the missed information could be searched from the Web or obtained based on existing information, the images can solve the query. If not, the images cannot solve the query.>, "correct": <According to the above reasoning, if you consider the images reasonable for the query to be solved by the tools, set the value to ’yes’, otherwise set the value to ’no’.>, "updated_query": <If you judge the correctness as ’no’, please rewrite the query to make it more relevant to the given images. If you judge the correctness as ’yes’, please output "no revision is needed." > } ”’
Figure 11: System prompt for the query-file verification.
Following are images, the query: <query>, inference whether the images can solve the query based on the perception ability, reasoning ability, and information search ability of an AI agent.
Figure 12: User prompt for the query-file verification.
F.4 Prompt for Step Verifier
@@
To evaluate the quality of intermediate steps taken by an agent during task execution, we design a step verifier consisting of two key components: a system prompt and a user prompt. The system prompt (Figure 13) provides the verifier model with detailed instructions for evaluating multiple candidate steps (‘CURRENT_STEP’) based on their coherence, logical progression, and effectiveness in advancing the task. It guides the model to consider contextual alignment with the prior step, tool usage, hallucination, and content relevance. The verifier is required to select the best step and justify its decision in a structured json format.
@@
The following are the given task, results of previous steps, and result of the current step. TASK: <task>
Step Sets: <step_set> # Format of step_set:
#     ‘PREVIOUS_RESULT‘: <The prior results obtained by the agent.>
#     ‘CURRENT_STEP‘: <The agent’s output, containing a ‘thought‘ and ‘code‘ intended to complete the task based on the observation.>
#     ‘CURRENT_RESULT‘: <The result or state produced by executing the ‘CURRENT_STEP‘.>
Now, you need to determine the best of the current steps based on the above information.
@@
Appendix G User Study Interface
G.1 Preference Alignment Study
Figure˜15(a) presents the web interface used to evaluate how well our automated verifier’s preferences align with those of human judges. In each trial, participants were shown a single task case along with a collection of candidate next‐step actions (each consisting of a brief “Thought” description and an optional code snippet or tool invocation). These options were the same ones ranked by our verifier, but presented in random order to prevent positional bias.
@@ G.1 Preference Alignment Study
By comparing the human‐selected option against the top choice of the verifier, we compute an agreement rate for each model and task type. High agreement indicates that the verifier captures human judgments of step quality; lower agreement reveals areas where the verifier’s ranking diverges from human intuition. These results were aggregated over 50 cases per participant, enabling both per‐case analysis and overall statistics on human–verifier alignment.
G.2 Data quality
Figure˜15(b) illustrates the web interface employed in our user study. For each case, participants proceeded through two consecutive scoring phases:
@@ G.2 Data quality
(a) Interface for the user study on verifier performance. Users select their preferred steps.
(b) Interface for the user study on task execution and trajectory. Users evaluate the quality of tasks and trajectories.
@@
Appendix H More Visualization
To further demonstrate the effectiveness and interpretability of our data generation and selection strategy, we present additional case studies in  Figure˜17,  Figure˜18, and  Figure˜16. These visualizations provide qualitative insights into how our method synthesizes diverse and informative training instances, and how the selection mechanism prioritizes samples that contribute meaningfully to model learning.
Figure 16: Case study for data generation and selection.
Figure 17: Case study for data generation and selection.
Figure 18: Case study for data generation and selection.